3) Assuming that firms compete a la Cournot, that all firms have the same marginal cost, and

that demand is linear, when is price most sensitive to changes in marginal cost: in a market

with very few firms or in a market with many firms? Show this formally. [Hint: assume

demand 𝑝 = π‘Ž – 𝑏𝑄]

b) Consider a Bertrand duopoly with differentiated products. Demand curves are given by

𝑝! = 600 βˆ’ 2π‘ž! βˆ’ π‘ž"

𝑝" = 600 βˆ’ π‘ž! βˆ’ 2π‘ž"

Suppose that the cost functions are given by (π‘ž#) = 60π‘ž# , for 𝑖 = 1, 2. Find the equilibrium

outputs, the prices and the profits.


1
Expert's answer
2021-10-18T11:28:05-0400

a.

Price is most when demand is elastic. It can be expressed as p=a-bQ meaning consumers will buy consume more when price is less .

b.

"\frac{dTC}{Q}=\frac{60Q}{Q}=60"


Price ="600-60-120=420"

Profit"P\u00d7Q=420\u00d760=25,200"